OpenStudy (anonymous):

a freight train and an express train leave towns 390 km apart, traveling toward one another. The freight train travels 30 km per hr slower than the express train. They pass one another 3 hr later. What are their speeds? use a system of equations

OpenStudy (anonymous):

One thing we do know is that the total distance traveled by the two trains = 390km : Let s = speed of the freight then (s+30) = speed of the express : Write a distance equation: Dist = time * speed : Freight dist + Express dist = 390 km 3s + 3(s+30) = 390 3s + 3s + 90 = 390 6s = 390 - 90 6s = 300 s = 300%2F6 s = 50 km/hr speed of the freight and 50 + 30 = 80 km/hr speed of the express : : Check solution by find the total dist 3(50) + 3(80) = 390

OpenStudy (anonymous):

I see where i messed up now thanks. Oh how would i rite this in a sentence?

OpenStudy (anonymous):

and is this a system of equation format? @andrea408

OpenStudy (anonymous):

@dpaInc